# A. CamillettiAI weather from Euro-Atlantic regimes

### v1

@@ 5.2 Impact of the number of WR on the reconstruction
In Fig. 4, the reconstruction skills (MAE, ACC, and CE) of the AI-model for monthly mean two-meter temperature and total precipitation anomalies during winter (DJF) and summer (JJA) are compared across configurations using seven WR indices (red), four WR indices (orange), the NAO index (light blue), and no indices (blue) as input. Analogously, Tab. 1 report the medians of the same metrics and the uncertainties on the last significant digit (reported in parenthesis), computed as the semi-difference of the skill metrics’ medians across multiple model initializations with different random seeds. The model trained without indices (No index) consistently shows the lowest skill scores for both temperature and precipitation. Despite this, it retains limited skill for two-meter temperature, with median ACC around $0.30\pm 0.07$ in winter and $0.32\pm 0.03$ in summer. This likely results from the AI-model capturing the long-term climate trend through the year and month inputs. However, the spatial distribution of this correlation lacks statistical significance in large regions, as shown in Fig. C.1. The inclusion of the NAO index, leads to some improvements. Since the NAO alone explains approximately 25% (16 %) of the atmospheric circulation variance in winter (summer), it significantly enhances the ACC for the precipitation in winter. Interestingly, for temperature, the skills are comparable between the models using the NAO index and without indices, indicating that much of the reconstruction skill for temperature likely arises from seasonal and anthropogenic trends rather than synoptic variability. Training the model with four WR indices yields substantial gains over the NAO model across both seasons and variables. For temperature, the typical improvements in ACC are about 29% in winter and 18% in summer, while for precipitation, they reach 100% (DJF) and 220% (JJA). The CE also improves, particularly for winter precipitation. Although CE remains near zero or slightly negative in JJA, particularly for precipitation, the four-index model exhibits far more dynamical skill and robustness than models with fewer indices. Expanding to seven WR indices results in smaller yet consistent improvements. The MAE remains largely comparable to the four WR setup, while ACC increases by 39% (DJF) and 20% (JJA) for temperature and by 15% (DJF) and 19% (JJA) for precipitation. The CE improves by approximately 260% for temperature and 41% for precipitation in winter but shows no substantial gains in summer, where skill levels remain low.
@@
The gains between four and seven indices are strongly dependent on the season. While we do not observe substantial improvements in summer, using seven WR gives considerable enhancement in winter. We conclude that using at least four WR indices can provide an effective balance between reconstruction performance and index predictability, offering robust and meaningful reconstruction skills for both temperature and precipitation in winter and summer.
@@
MAE ACC CE
7 WR 4 WR NAO No idx 7 WR 4 WR NAO No idx 7 WR 4 WR NAO No idx
DJF 0.29(1) 0.35(2) 0.36(1) 0.36 0.61(4) 0.44(8) 0.34(4) 0.30(7) 0.29(4) 0.08(6) 0.02(1) -0.06(3)
JJA 0.46(2) 0.46(1) 0.49(2) 0.54(2) 0.48(4) 0.40(4) 0.34(2) 0.32(3) 0.05(4) 0.03(2) -0.06(5) -0.17(4)
@@ 7 WR 4 WR NAO No idx 7 WR 4 WR NAO No idx 7 WR 4 WR NAO No idx
Table 1: Median of MAE, ACC, and CE for reconstructed monthly mean two-meter temperature and total precipitation during winter (DJF) and summer (JJA), using seven WR, four 4 WR, NAO, or no indices as input to the AI model. Uncertainties on the last significant digit (in parentheses) denote the semi-difference (half the difference between maximum and minimum) of the skill metrics’ medians across multiple model initializations with different random seeds. Uncertainties are not reported when different initialization seeds give consistent results.
5.3 Reconstruction degradation with indices errors
@@
The 50 perturbed indexes for each value of the perturbation MARE are used as input to the AI model. The MAE, the ACC, and the CE for the summer and winter seasons can be evaluated as usual for each of the 50 reconstructed variables. After the skills are computed for each realization, we perform an average over the random realizations to evaluate the skills for the given MARE.
Figure 5: Spatial median of the mean two-meter temperature (red) and total precipitation (blue) ACC (triangles) and CE (rhomboids) reconstructed by the AI model in the winter seasons (top panel) and summers (bottom panel) from 2011 to 2024 , as a function of index MARE. The values are averaged over 50 random realizations of $I_{\rm wr}^{\prime}$ sampled from $I_{\rm wr}\times N\left(0,\sqrt{\pi/2}~{}{\rm{MARE}}\right)$. Horizontal lines are the SEAS5 spatial median of the ACC (solid) and CE (dashed) for the same variables and seasons.
@@
The spatial median of the mean two-meter temperature and total precipitation ACC and CE, averaged over the 50 random realizations of the perturbed indices, are shown in Fig. 5 for the 2011 – 2024 winters and summers. Increasing the MARE on the indices, the AI model skills naturally degrade up to a point where they can be worse than the SEAS5 ensemble-mean ACC and CE. Both in winter and summer, the median ACC of the reconstructed mean two-meter temperature and total precipitation are always better than the SEAS5 ensemble-mean forecast ACC, indicating a strong efficacy of the AI model in capture the correlation between the Euro-Atlantic WR and European two-meter temperature and precipitation also in the presence of Gaussian distributed errors on the input indices. As for the ACC, the median CE of the mean two-meter temperature in winter and the precipitation in summer is always higher than the respective SEAS5 ensemble-mean median CE. More interesting is the case of the median CE of the reconstructed total precipitation in winter and mean two-meter temperature in summer, which becomes worse than the relative SEAS5 skill at MARE $\sim 80\%$ and $120\%$, respectively. Taking the lowest value of MARE of 80 % as a reference, this poses a lower bound on the forecast accuracy of the seven $I_{\rm wr}$ required to have a seasonal forecast of the mean two-meter temperature and total precipitation with higher deterministic skills than the bias-corrected SEAS5 ensemble mean forecast used in this work. We can express the results obtained here in terms of MAE of the perturbed indexes, $\langle|I_{\rm wr}-I_{\rm wr}^{\prime}|\rangle=0.35$ for MARE of 80 %.
5.4 AI-model in a concrete forecasting scenario
@@
To test the AI model in a real forecasting scenario, we now analyze the reconstruction of the monthly mean two-meter temperature and total precipitation when using the monthly mean seven WR indices forecasted by the bias-corrected SEAS5, $I_{\rm wr}^{\rm SEAS5}$ (see Sec. 2.2), as model inputs. For both the AI model and SEAS5 the skills are computed for the ensemble mean, where the ensemble of the AI model reconstruction is obtained using as input the seven $I_{\rm wr}^{\rm SEAS5}$ forecast of each ensemble member. In the following, the SEAS5 forecasts for the winters (summers) are initialized on the first of November (May) to predict December, January and February (June, July and August).
@@
Figure 6: MAE (top row), ACC (middle row) and CE (bottom row) for the SEAS5 (red) and the AI model (blue) using $I_{\rm wr}^{\rm SEAS5}$ as input. The first and second columns show the winter (left half) and summer (right half) skills for the two-meter temperature and precipitation, respectively.
@@
Fig. 6 presents a comparative assessment of the hybrid NWP-AI model and the SEAS5 forecasting system for boreal winter and summer, evaluated using the usual metrics. The NWP-AI model typically exhibits better or comparable skills to the SEAS5 forecast. In particular, it achieves lower MAE values, particularly for the winter mean two-meter temperature and summer total precipitation anomalies. This enhanced performance is likely attributable to the use of a MSE loss during training, which suppresses variability in the model output. In contrast, SEAS5 tends to produce forecasts with higher variability, often resulting in higher MAE. The reduced MAE of the NWP-AI model translates into higher values of the CE, especially in cases where the discrepancy in MAE relative to SEAS5 is substantial. Nevertheless, both SEAS5 and NWP-AI model yield CE values that are often close to or below zero, indicating limited skill relative to the climatology of the test dataset. Regarding the ACC, the NWP-AI model slightly outperforms SEAS5 for the two-meter temperature in both seasons, but the ACC are comparable for the total precipitation anomalies.
@@ 6 Conclusions
This study presents and evaluates a lightweight AI-based model for reconstructing seasonal anomalies of two-meter temperature and total precipitation in Europe using Euro-Atlantic WR indices. The model demonstrates robust skill in capturing both spatial and temporal variability of temperature, with somewhat reduced performance for precipitation. Reconstruction skills are higher during winter, reflecting the stronger influence of large-scale circulation patterns on surface climate during this season. We test the model’s performance by varying the number of WR indices used as input. Despite the best skills being naturally reached by the models using seven WR, four indices represent a reasonable trade-off between reconstruction accuracy and forecast feasibility. We further assess the model’s robustness to errors in the input WR indices and identify a lower bound on their forecast accuracy necessary to outperform a state-of-the-art dynamical forecasting system, SEAS5. Specifically, to exceed SEAS5’s skill, the indices of winter (DJF) and summer (JJA) must be predicted with a MARE below approximately 80%, corresponding to a MAE of about 0.35. Finally, in a realistic forecasting setup where the WR indices are derived from bias-corrected SEAS5 forecasts, the hybrid NWP-AI model achieves comparable or superior performance relative to SEAS5, particularly for winter temperature and summer precipitation. These results highlight the potential of AI-driven approaches, grounded in dynamical predictors like WR indices, to complement traditional numerical forecasting methods for improved long-range climate prediction.
Nonetheless, several limitations remain. Training AI models usually require a large amount of data. We use monthly values from 1940 to 2010 , for a total of 852 points. The small training dataset forces the use of a lightweight model with a limited number of trainable parameters. Additionally, training with the MSE loss promotes smooth reconstructions with reduced variability. Other losses could be tested to have more realistic reconstructions. Moreover, the model is tested during the winters and summers in the 2011 – 2024 years. A small test dataset can significantly influence the statistical stability of the results presented in this work.
@@
Despite these limitations, the approach is flexible and adaptable. The model can be extended to other regions, time scales, and surface variables. In particular, the sub-seasonal time scale presents a promising opportunity, offering both a richer dataset and potential for operational forecasting improvements. Recent studies have proposed new approaches for computing the WR [Paepe and Cruz, 2025, Spuler et al., 2024, Springer et al., 2024, Mukhin et al., 2024a]. Incorporating these WR could potentially enhance the reconstruction performance of the proposed AI-model.
@@
The next efforts will be directed towards a fully AI-driven method to forecast the WR indices, which can be used as input to the AI model presented here to obtain a forecast of the surface-level variables.
[…]
@@
The authors thank Ilaria Tessari and Federico Fabiano for the useful discussions on the year-round WR calculations. We also thank Leonardo Aragao for the confrontation on the bias-correction of SEAS5. This work was developed with financial support from ICSC–Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, Spoke4 - Earth and Climate, funded by European Union – NextGenerationEU.
@@
Daily post-processed ERA5 data on surface [Hersbach et al., 2024b] and pressure levels [Hersbach et al., 2023], was generated using Copernicus Climate Change Service information 2020, obtained from the Copernicus Climate Change Service [Hersbach et al., 2024a, b]. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The AI-model implemented in this work are written in PyTorch Lightning [Falcon et al., 2020] and released under the MIT License through a GitHub repository (in preparation) based on the Lightning-Hydra-Template [Yadan, 2019]. The repository also contains all the script necessary to reproduce the results discussed in the paper (figures and tables). The data used to train the models will be made available on Zenodo (in preparation).

### v2

@@ 5.3 Impact of the number of WR on the reconstruction
Figure 6: 2011 – 2024 winters (first and third columns) and summers (second and fourth columns) MAE (top row), ACC (middle row) and CE (bottom row) of the two-meter temperature anomalies and precipitation anomalies, reconstructed by the AI models trained with seven WR (red), four WR (orange), NAO (light-blue) and zero (blue) indices. The box plots summarize the spatial data, showing the median (central line), interquartile range (box), and range.
@@
In Fig. 6, the reconstruction skills (MAE, ACC, and CE) of the AI models for monthly mean two-meter temperature and total precipitation anomalies during winter (DJF) and summer (JJA) are compared across configurations using seven WR indices (red), four WR indices (orange), the NAO index (light blue), and no indices (blue) as input. Analogously, Tab. 1 and Tab. 2 report the medians of the same metrics and the uncertainties on the last significant digit reported in parentheses. The uncertainties are computed as the semi-difference of the skill metrics’ medians across multiple model initializations with different random seeds, to quantify the spread due to the random seed initialization. The models trained without indices (No index) consistently show the lowest scores for both temperature and precipitation. Despite this, the model retains limited skill for two-meter temperature, with median ACC around $0.28\pm 0.03$ in winter and $0.32\pm 0.03$ in summer. This likely results from the AI model capturing the long-term climate trend through the year and month inputs. However, the spatial distribution of this correlation lacks statistical significance in large regions, as shown in Fig. D.1. The inclusion of the NAO index leads to some improvements. Since the NAO alone explains approximately 25% (16 %) of the atmospheric circulation variance in winter (summer), it significantly enhances the ACC for both variables in winter. Interestingly, for temperature in summer, the skills are comparable between the models using the NAO index and without indices. We speculate that much of the reconstruction skill for the summer temperature likely arises from anthropogenic trends, with only a comparatively small contribution from the NAO. A thorough assessment of this hypothesis is beyond the scope of the present work. The reconstruction of summer precipitation has instead a general improvement with respect to the model trained without indices. Training the models with four WR indices yields substantial gains over the NAO models across both seasons and variables. For temperature, the typical improvements in ACC are about 28% in winter and 16% in summer, while for precipitation, they reach 74% (DJF) and 100% (JJA). The CE also improves, particularly for the winter season. Although CE remains near zero or slightly negative in summer, the four-index models exhibit far more dynamical skill and robustness than models with fewer indices. Expanding to seven WR indices results in reduced yet consistent improvements. The MAE remains largely comparable to the four WR setup, while ACC increases by 27% (DJF) and 35% (JJA) for temperature and by 17% (DJF) and 14% (JJA) for precipitation. The CE improves by approximately 113% for temperature and 45% for precipitation in winter, but remains low in summer.
@@
In view of a data-driven forecast of WR indices followed by the anomalies reconstruction, we suggest that using at least four WR indices can provide an effective balance between reconstruction performance and index predictability, offering robust and meaningful reconstruction skills for both temperature and precipitation in winter and summer.
@@
Season Metric 7 WR 4 WR NAO No idx
DJF MAE (K) 0.78(3) 0.91(5) 1.04(6) 1.01(2)
ACC 0.65(4) 0.51(1) 0.40(2) 0.28(3)
CE 0.32(7) 0.15(2) -0.01(4) -0.04(1)
JJA MAE (K) 0.77(2) 0.88(6) 0.93(3) 0.82(2)
ACC 0.50(5) 0.37(3) 0.32(1) 0.32(3)
CE 0.07(5) -0.03(6) -0.09(3) -0.08(2)
Table 1: Median MAE, ACC, and CE for monthly mean two-meter temperature in DJF and JJA using seven WR, four WR, NAO, or no index as AI model input. Uncertainties on the last significant digit (in parentheses) denote the semi-difference (half the difference between maximum and minimum) of the skill metrics’ medians across multiple model initializations with different random seeds.
@@
Table 2: As Tab. 1 but for monthly mean total precipitation in DJF and JJA using seven WR, four WR, NAO, or no index as AI model input. Uncertainties on the last significant digit (in parentheses) denote the semi-difference (half the difference between maximum and minimum) of the skill metrics’ medians across multiple model initializations with different random seeds.
5.4 Reconstruction degradation with index errors
@@
The 50 perturbed indices for each value of the perturbation MARE are used as input to the AI models. The MAE, the ACC, and the CE for the summer and winter seasons can be evaluated as usual for each of the 50 reconstructed variables. After the skills are computed for each realization, we perform an average over the random realizations to evaluate the skills for the given MARE.
@@
The spatial median of the mean two-meter temperature and total precipitation ACC and CE, averaged over the 50 random realizations of the perturbed indices, are shown in Fig. 7 for the 2011 – 2024 winters and summers. Increasing the MARE of the perturbed indices, the AI models’ skills naturally degrade up to a point where they can be worse than the SEAS5 ensemble-mean ACC and CE. Both in winter and summer, the median ACC of the reconstructed mean two-meter temperature and total precipitation is always better than the SEAS5 ensemble-mean forecast ACC, indicating a strong efficacy of the AI models in capture the correlation between the Euro-Atlantic WR and European two-meter temperature and precipitation also in the presence of Gaussian distributed errors on the input indices. As for the ACC, the median CE of the mean two-meter temperature in winter and the precipitation in summer is always higher than the respective SEAS5 ensemble-mean median CE. More interesting is the case of the median CE of the reconstructed total precipitation and mean two-meter temperature, which becomes comparable to the relative SEAS5 skill at MARE = 100%. Taking the lower value of MARE of 80% as a reference, this poses a lower bound on the forecast accuracy of the seven $I_{\rm wr}$ required to have a seasonal forecast of the mean two-meter temperature and total precipitation with higher deterministic skills than the bias-corrected SEAS5 ensemble mean forecast used in this work. We can express the results obtained here in terms of MAE of the perturbed indices, $\langle|I_{\rm wr}-I_{\rm wr}^{\prime}|\rangle=0.35$ for MARE of 80%.
Figure 7: Spatial median of the mean two-meter temperature (red) and total precipitation (blue) ACC (triangles) and CE (rhomboids) reconstructed by the AI models in the winter seasons (top panel) and summers (bottom panel) from 2011 to 2024 , as a function of index MARE. The values are averaged over 50 random realizations of $I_{\rm wr}^{\prime}$ sampled from $I_{\rm wr}\times N\left(0,\sqrt{\pi/2}\penalty 10000\ {\rm{MARE}}\right)$. Horizontal lines are the SEAS5 spatial median of the ACC (solid) and CE (dashed) for the same variables and seasons.
5.5 AI-models in a concrete forecasting scenario
@@
To test the AI models in a real forecasting scenario, we now analyze the reconstruction of the monthly mean two-meter temperature and total precipitation when using the monthly mean seven WR indices forecasted by the bias-corrected SEAS5, $I_{\rm wr}^{\rm SEAS5}$ (see Sec. 2.2), as model inputs. Although the AI models are not inherently designed for ensemble forecasting, since they are trained deterministically using the MSE loss, an ensemble can nevertheless be constructed by driving the models with the seven $I_{\rm wr}^{\rm SEAS5}$ forecasts from each SEAS5 ensemble member. This allows us to evaluate not only deterministic metrics based on the ensemble means, but also probabilistic scores such as the CRPS and the SSR for both the hybrid NWP-AI models and SEAS5. In the following, the SEAS5 forecasts for the winters (summers) are initialized on the first of November (May) to predict December, January, and February (June, July, and August).
@@
Figure 8: CRPS (first row), SSR (second row), ACC (third row) and CE (last row) for the SEAS5 (red) and the AI models using $I_{\rm wr}^{\rm SEAS5}$ as input (blue). The first and second columns show the winter (left half) and summer (right half) skills for the two-meter temperature and precipitation, respectively. The box plots summarize the spatial data, showing the median (central line), interquartile range (box), and range.
@@
Fig. 8 presents a comparative assessment of the hybrid NWP-AI models and the SEAS5 forecasting system for boreal winter and summer. The NWP-AI models typically exhibit better or comparable skills to the SEAS5 forecast. In particular, they achieve lower CRPS values, particularly for the winter mean two-meter temperature and summer total precipitation anomalies. This enhanced performance is likely attributable to the use of an MSE loss during training, which suppresses variability in the models’ output. In contrast, SEAS5 tends to produce forecasts with higher variability, often resulting in higher MAE but with an SSR near the unity. The overconfidence of the hybrid NWP-AI approach is its major drawback: the ensemble spread is markedly reduced by the AI model, which is not adequately designed and trained to face a probabilistic forecast. In the conclusions, we outline possible directions for addressing this limitation. The reduced MAE of the NWP-AI models translates into higher values of the CE, especially in cases where the discrepancy in MAE relative to SEAS5 is substantial. Nevertheless, both SEAS5 and NWP-AI models yield CE values that are often close to or below zero, indicating limited skill relative to the climatology of the test dataset. Regarding the ACC, the NWP-AI models slightly outperform SEAS5 for the two-meter temperature in both seasons. For the total precipitation anomalies, the ACC is comparable in winter and worse in summer.
Considering the stability of the AT-model in the presence of Gaussian-distributed errors in the WR indices, as discussed in Sec. 5.4, the relatively poor performance of the NWP-AI models compared to the AI model driven by perfect indices indicates that the SEAS5 forecast of the WR (i.e., $Z500$) has limited skill. Indeed, the CRPS of the indices degrades rapidly with forecast lead time, reaching around 0.6 already for forecasts of the month following the initialization (see Appendix E).
@@ 6 Conclusions
This study presents and evaluates two lightweight AI based models for reconstructing monthly anomalies of two-meter temperature and total precipitation in Europe using Euro-Atlantic WR indices. The models demonstrate robust skill in capturing both spatial and temporal variability of temperature, with somewhat reduced performance for precipitation. Reconstruction skills are higher during winter, reflecting the stronger influence of large-scale circulation patterns on surface climate during this season. We test the performance of the models by varying the number of WR indices used as input. Despite the models using seven WR achieving the best skills, four indices represent a reasonable trade-off between reconstruction accuracy and forecast feasibility. We further assess the models’ robustness to errors in the input WR indices and identify a lower bound on their forecast accuracy necessary to outperform a state-of-the-art dynamical forecasting system, SEAS5. Specifically, to exceed SEAS5’s skill, the indices of winter (DJF) and summer (JJA) must be predicted with an MARE below approximately 80%, corresponding to a MAE of about 0.35 when the indices are normalized by the mean and standard deviation of the 1981 - 2011 climatology. Finally, in a realistic forecasting setup where the WR indices are derived from bias-corrected SEAS5 forecasts, the hybrid NWP-AI models achieve comparable or superior performance relative to SEAS5, particularly for winter temperature and summer precipitation. These results highlight the potential of AI-driven approaches, grounded in dynamical predictors like WR indices, to complement traditional numerical forecasting methods for improved long-range climate prediction.
Nonetheless, several limitations remain. Training AI models usually requires a large amount of data. We use monthly values from 1940 to 2005 , totaling 852 points. The small training dataset forces the use of a lightweight architecture with a limited number of trainable parameters. Training with the MSE loss encourages smooth reconstructions with reduced variability. While this leads to lower error scores, it also suppresses part of the natural variability of the climate system. This behavior is typical of a deterministic model trained on pointwise losses, which tend to regress toward the mean state rather than reproduce the full range of plausible outcomes. A natural extension would be to design the architecture for probabilistic reconstruction. For instance, an ensemble of reconstructions could be generated by training the network with proper probabilistic losses such as the CRPS. This would allow the model to produce a distribution of possible realizations rather than a single deterministic trajectory, better capturing forecast uncertainty and enabling the use of ensemble-based verification metrics.
@@
Despite these limitations, the approach is flexible and adaptable. The model can be extended to other regions, time scales, and surface variables. In particular, the sub-seasonal time scale presents a promising opportunity, offering both a richer dataset and potential for operational forecasting improvements. Recent studies have proposed new approaches for computing the WR [geert_2025, spuler_identifying_2024, springer_unsupervised_2024, mukhin_revealing_2024]. Incorporating these WR could potentially enhance the reconstruction performance of the proposed AI model.
@@
The next efforts will be directed towards a fully probabilistic AI-driven method to forecast the WR indices, which can be used as input to the AI model presented here (or to a probabilistic extension) to obtain a forecast of the surface-level variables.
Acknowledgements
@@
The authors thank Ilaria Tessari and Federico Fabiano for the valuable discussions on the year-round WR calculations. We also thank Leonardo Aragao for the confrontation on the bias-correction of SEAS5. This work was developed with financial support from ICSC–Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, Spoke4 - Earth and Climate, funded by European Union – NextGenerationEU.
@@
Daily post-processed ERA5 data on surface [ERA5Copernicus_surface_daily] and pressure levels [ERA5pressure_daily], was generated using Copernicus Climate Change Service information 2020, obtained from the Copernicus Climate Change Service [ERA5Copernicus_pressure_daily, ERA5Copernicus_surface_daily]. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The AI models implemented in this work are written in PyTorch Lightning [william_falcon_2020_3828935] and released under the MIT License through a GitHub repository (https://github.com/DSIP-FBK/anom_recon) based on the Lightning-Hydra-Template [Yadan2019Hydra]. The repository also contains all the scripts necessary to reproduce the results discussed in the paper (figures and tables). The data used to train the models is available on Zenodo (https://zenodo.org/records/17276947).
Author contributions
Alessandro Camilletti: Investigation, Conceptualization, Methodology, Data Curation, Formal Analysis, Software, Validation, Writing – Original Draft. Gabriele Franch: Methodology, Software, Writing – Review & Editing, Resources, Supervision. Elena Tomasi: Supervision, Writing – Review & Editing. Marco Cristoforetti: Funding acquisition, Project administration, Writing – Review & Editing, Supervision. All authors participated in discussions throughout the development of the study and contributed to refining the manuscript. All authors have read and approved the final version of the manuscript.
Conflict of interest
All authors declare that they have no conflicts of interest.